The company's competitors we afraid because the company's production increased exponentially. Only five years after the company began its production, the prodiction was 3 times greater than when production began. How many times greater will the production be at the end of 15 years?
Answer by greenestamps(13200)



My answer is probably not the answer the author of the problem wanted, because the problem is very probably not written the way the author intended.

The trouble with the statement of the problem is that "3 times GREATER THAN x" means x, plus 3 times x, which makes 4x. So "3 times greater than" means "4 times as great".

Unfortunately, in sloppy everyday English usage, "3 times greater than" and "3 times as great as" are usually used to mean the same thing.

Grammatically, the correct interpretation of the given information is that the company's production increases by a factor of 4 every 5 years.

That means in 15 years (3 5-year periods) the production will increase by a factor of 4^3 = 4*4*4 = 64.

And that means the company's production at the end of 15 years will be 63 times GREATER THAN it is now.

CORRECT ANSWER: 63

---------------------
WRONG ANSWER: 27. (The company's production grows by a factor of 3 every 5 years, so in 15 years it will grow by a factor of 3*3*3=27.)